### v1

# LLM Test Generation via Iterative Hybrid Program Analysis

Abstract.
Automating unit test generation remains a significant challenge, particularly for complex methods in real-world projects. While Large Language Models (LLMs) have made strides in code generation, they struggle to achieve high branch coverage due to their limited ability to reason about intricate control flow structures. To address this limitation, we introduce Panta, a technique that emulates the iterative process human developers follow when analyzing code and constructing test cases. Panta integrates static control flow analysis and dynamic code coverage analysis to systematically guide LLMs in identifying uncovered execution paths and generating better test cases. By incorporating an iterative feedback-driven mechanism, our technique continuously refines test generation based on static and dynamic path coverage insights, ensuring more comprehensive and effective testing. Our empirical evaluation, conducted on classes with high cyclomatic complexity from open-source projects, demonstrates that Panta achieves 26% higher line coverage and 23% higher branch coverage compared to the state-of-the-art.
Test generation, program analysis, control-flow analysis, large language models
1. Introduction
Automated test generation plays a crucial role in software engineering by enhancing code reliability and maintainability. Large Language Models (LLMs) can replicate human coding styles and generate test cases with meaningful variable names, leveraging patterns learned from extensive training data. This capability has led to the adoption of LLMs for test code generation (Nashid et al., 2023; Lemieux et al., 2023; Alshahwan et al., 2024; Chen et al., 2024; Pan et al., 2025; Ryan et al., 2024; Wang et al., 2024a; Deljouyi et al., 2024).
Despite their effectiveness in general coding tasks, LLMs—trained on diverse datasets—often struggle with the unique code structures and logic of specific projects. This limitation is particularly evident in cases involving complex control flow and conditional statements. Recent studies (Siddiq et al., 2024; Schäfer et al., 2023; El Haji, 2023; Wang et al., 2024b) have shown that LLM-generated test cases frequently fail to achieve sufficient branch coverage, as they inadequately capture intricate conditional logic.
To address this limitation, researchers have explored techniques that supplement LLMs with additional program-related information, such as focal method slices (Wang et al., 2024a) and path constraints (Ryan et al., 2024). These approaches rely on static prompts containing method fragments to generate tests for each independent segment. However, they lack mechanisms to guide the generation process toward areas requiring additional coverage. Furthermore, since LLMs do not always produce correct code, only a subset of the generated tests compile and pass, leaving significant portions of the code untested. Additionally, these methods primarily focus on generating new test classes rather than augmenting existing ones. CoverUp (Pizzorno and Berger, 2024) takes a different approach by leveraging code coverage information to highlight uncovered portions of the code. However, our empirical evidence suggests that this approach is insufficient for more complex code, where additional contextual guidance is needed to drive meaningful improvements in branch coverage.
(a) Execute existing tests to locate code that lacks coverage
1 public static Options parsePattern (String pattern){
2 char opt = ;
3 boolean required = false;
4 Class<?> type = null;
5 final Options options = new Options();
6 for (int i = 0; i < pattern.length(); i++){
7 final char ch = pattern.charAt(i);
8 if (!isValueCode(ch)) {
9 // omit some details
10 opt = ch;
11  } else if (ch == ’!’) {
12  required = true;
13 } else {
14 type = (Class<?>) getValueClass(ch);
15 }
16 }
17  if (opt != ) {
18 // omit some details
19 }
20 return options;
21 }
(b) Identify paths and reason about execution scenarios
2 char opt = ;
3 boolean required = false;
4 Class<?> type = null;
5 final Options options = new Options();
6 for(int i=0;i<pattern.length();i++)is False
17  if (opt != ) is False
20 return options;
When pattern is empty
2 char opt = ;
3 boolean required = false;
4 Class<?> type = null;
5 final Options options = new Options();
6 for(int i=0;i<pattern.length();i++)is True
7 final char ch = pattern.charAt(i);
8 if (!isValueCode(ch)) is False
11  else if (ch == ’!’) is True
12  required = true;
6 for(int i=0;i<pattern.length();i++)is False
17  if (opt != ) is False
20 return options;
When pattern contains only character ”!”
(c) Add new tests for the execution scenarios
@Test
public void testEmptyPattern() {
Options options =
PatternOptionBuilder.parsePattern("");
assertTrue(options.getOptions().isEmpty());
@Test
public void testRequiredOption() {
Options options =
PatternOptionBuilder.parsePattern("!");
// proper assertions
Figure 1. Workflow of a developer to add new test cases
In this paper, we present Panta (Program ANalysis guided Test Automation), a novel approach that combines dynamic code coverage analysis and static control flow analysis to generate tests for under-tested execution paths. Our approach follows an iterative workflow, implementing a multi-step, feedback-driven fully automated process. Panta first gathers dynamic code coverage information and static control flow insights. Based on the extracted program analysis information, it identifies and selects execution paths that require additional test coverage for the class, and prompts an LLM to generate test cases that cover those paths. After test generation, Panta validates the generated tests, adding the passing ones back to the test class while repairing failed tests. The feedback from the test repair process is incorporated into subsequent test generation iterations.
Our work makes the following contributions:
  • An iterative feedback-driven technique that integrates dynamic and static program analysis to assist LLMs in generating tests that cover complex control flow paths.
  • A novel path selection strategy that first ranks linearly independent execution paths—identified via control flow analysis—using a coverage deficiency score computed from code coverage data. It then selects prioritized paths while incorporating historical selection data to effectively balance exploitation and exploration.
  • A fully automated test generation tool, Panta, that supports both generating new test classes and augmenting existing tests.
  • An empirical evaluation of our approach using real-world open-source programs from the Defects4J (Just et al., 2014) benchmark. Our study demonstrates that Panta achieves 26% higher line coverage and 23% higher branch coverage compared to the state of the art.
2. Motivation
Recent advances in LLMs have dramatically enhanced their code generation capabilities, including the ability to create tests for diverse execution scenarios. However, when dealing with complex methods that contain multiple conditional branches, LLMs often struggle to achieve comprehensive branch coverage.
As a running example, Figure 1, derived from the Cli project in the Defects4J dataset (Just et al., 2014), illustrates this challenge. We prompted Llama 3.3 70b (Llama, 2025), Meta’s latest open-source model, to generate tests for the parsePattern method shown in Figure 1 and then measured the resulting coverage. The generated tests, however, failed to cover the true branch of the else if (ch == ‘!’) condition at Lines 1112 and the false branch of the if (opt != ’ ’) condition at Line 17. We also tried with GPT-4o, which similarly missed the false branch of the condition at Line 17.
Our insight is that emulating a developer’s test creation process can more effectively guide LLMs to generate test cases. Typically, developers begin by checking whether a corresponding test class exists. If it does not, they create a test class skeleton with the necessary imports and gradually add test cases. If it does exist, they run the tests to generate a coverage report that highlights untested statements. They then analyze the control flow—particularly conditional branches and loops—to identify missing execution paths, add new tests to cover these paths, and rerun the tests until adequate coverage is achieved.
For example, Figure 1 shows developers executing the tests and reviewing a JaCoCo (geeksforgeeks, 2025a) coverage report. In the parsePattern method, the report indicates that two branches remain uncovered. By analyzing the control flow, developers determine that reaching the true branch within a for loop requires the loop condition to be true while accessing the false branch outside the loop depends on the loop condition being false. Figure 1 illustrates two representative execution paths: one where an empty pattern prevents loop execution and one where a single ’!’ triggers the else if branch, which also covers the false branch. New test cases representing these scenarios are shown in Figure 1.
While developers iteratively refine and expand coverage until achieving the desired adequacy level, manually analyzing the code and creating tests remains a time-intensive task. To address this challenge, our work aims to automate and replicate the iterative workflow that developers follow when generating test cases by combining dynamic and static program analysis to effectively guide LLMs to cover hard-to-cover branches.
Figure 2. Overview of Panta.
3. Approach
Our approach, called Panta, leverages hybrid program analysis that seamlessly integrates static and dynamic analyses. By iteratively combining static control flow analysis with dynamic code coverage analysis, Panta enhances LLMs’ ability to determine where and how to generate test cases that improve branch coverage.
Figure 2 provides an overview of this process. It begins by scanning the entire project to identify directories containing source files and their corresponding test files. If a source file lacks an associated test file, a new test file skeleton is created in the designated test directory. For each source file, the hybrid approach first uses static control flow analysis to extract potential execution paths and then employs dynamic analysis—executing the corresponding test file—to collect real-time branch coverage information. The selected paths, derived from a unique path selection strategy discussed in Section 3.2, are incorporated into a test generation prompt that guides the LLM in generating tests.
In addition, the generated tests undergo validation: passing tests are integrated into the test file while failing tests are directed through a test repair procedure. If a test cannot be repaired, it is fed back into the test generation prompt as feedback for the next iteration. This iterative workflow continues until the stopping conditions defined in Section 3.4 are met. Since the process is applied uniformly across all source files and their corresponding test files, we detail the iterative test generation process for a single source file and its associated test file in the following subsections.
Input: srcFile
Output: methodDict
1 $CFG_{f}$ = CFGraph(srcFile)
2 MUTs = identifyMethodUnderTestFrom(srcFile)
3 /* extract paths for each method under test */
4 foreach method in MUTs do
5       $CFG_{m}$ = extractCFGraphForMethod($CFG_{f}$, method)
6       $CYC_{m}$= computeCyclomaticComplexity($CFG_{m}$)
7       $paths$ = ExplorePaths($CFG_{m}$)
8       methodDict[method] = ($CYC_{m}$, $paths$)
9
10 end foreach
11return methodDict
12
13/* Explore paths using BFS */
14 Function ExplorePaths($CFG_{m}$):
15       start, end = getStartEndNodeFrom($CFG_{m}$)
16       queue.append([start])
17       while queue is not empty do
18             for i = 0 to length(queue) - 1 do
19                   currentPath = queue.pop(0)
20                   lastNode = getLastNodeOf(currentPath)
21                   foreach successor in successors(lastNode) do
22                         nextPath = copy(currentPath)
23                         nextPath.append(successor)
24                         if successor is not the end node then
25                               queue.append(nextPath)
26
27                        else
28                               if nextPath contains edges not visited then
29                                     methodPaths.append(nextPath)
30
31                               end if
32                              if all edges of $CFG_{m}$ has been visited then
33                                     return methodPaths
34
35                               end if
36
37                         end if
38
39                   end foreach
40
41             end for
42
43       end while
44
45
46
Algorithm 1 Path Approximation through CFG
3.1. Static Path Approximation
Extracting all execution paths from a source file is often impractical due to the complexity introduced by loop conditions, which can lead to an exponential or unbounded number of paths. To address this, we propose a path approximation strategy through control flow analysis.
We introduce Algorithm 1, a path approximation method that utilizes the control flow graph (CFG) to extract a set of linearly independent paths, ensuring that each path introduces at least one new edge. We leverage an existing CFG generation technique (Das et al., 2023) to construct the control flow graph for the entire source file, denoted as $CFG_{f}$ (Line 1). Next, we identify the methods under test (MUTs), excluding helper methods such as private methods, getters, and setters (Line 1). For each MUT, we extract its corresponding method-level control flow graph ($CFG_{m}$) from $CFG_{f}$ (Line 1), forming the basis for our path approximation strategy.
Figure 3. Control flow graph for parsePattern in Figure 1.
Each $CFG_{m}$ is a fully directed graph where code blocks correspond to nodes, and control flow determines the directed edges. We compute the cyclomatic complexity (CYC) (McCabe, 1976) for each method (Line 1), defined as #$\#\text{edges}-\#\text{nodes}+2$. For instance, in Figure 3, we illustrate the $CFG_{m}$ for the parsePattern method from Figure 1. We use the line numbers from Figure 1 to represent nodes, where lines 2 to 5 and lines 9 and 10 form basic code blocks that are grouped into single nodes. Circles represent nodes without branches, while branch nodes are marked as pentagons. The directed edges indicate the control flow between nodes. To ensure a structured representation, we introduce two additional nodes: a start node (1:start) representing the method entry point and a dummy end node (21:end) to unify all execution paths into a single exit. These start and end nodes are depicted as octagons. Excluding these additional nodes, the parsePattern method consists of 12 nodes and 15 edges, with a cyclomatic complexity (CYC) value of 5. We then extract paths using the ExplorePaths in Algorithm 1 (Lines 11), which is based on breadth-first search (BFS). The algorithm collects the shortest set of paths, ensuring each includes at least one unique edge. Exploration terminates once all edges have been visited. By employing ExplorePaths, we guarantee that for each method under test, the number of extracted paths equals its CYC.
Using Algorithm 1, we obtain an approximated set of paths for each method under test in a static manner. Next, we use dynamic analysis to identify uncovered execution paths.
3.2. Dynamic Path Selection
We present a path selection strategy in Algorithm 2 that integrates dynamic code coverage data with the static paths extracted from Algorithm 1 to identify and prioritize uncovered paths for each test generation iteration.
Given that multiple test generation iterations may have already occurred, we maintain a pathHistory log to track path selection data from previous iterations. Since large language models (LLMs) do not always guarantee the generation of correct tests that fully cover the selected paths in each iteration, a path may be selected multiple times if it remains uncovered. However, if a path has been selected multiple times without being covered, it may indicate that the path itself is invalid or that LLMs struggle to reason about it, making further attempts unproductive. To optimize resource usage, we limit the number of selection attempts per path to maxSelectedConst. If a path has been selected maxSelectedConst times without being covered, it is excluded from future selections. The data variables, pathHistory and maxSelectedConst, are used globally throughout the entire test generation process.
Data: pathHistory, maxSelectedConst
Input: srcFile, testFile, methodDict
Output: selectedPaths
1
2// Run tests and collect coverage information
3 runTestsWithCoverage(testFile)
4 missedLines, missedBranches = parseCoverageReportFor(srcFile)
5 selectedPaths = dict()
6 // Process each method under test
7 foreach method in methodDict do
8       methodPaths = methodDict[method][1]
9       // Filter paths
10       candidatePaths = []
11       foreach path in methodPaths do
12             missedScore = count(missedLines $\cap$ path) + count(missedBranches $\cap$ path)
13             if missedScore ¿ 0 then
14                   selectedCount = pathHistory [method][path]
15                   if selectedCount ¡ maxSelectedConst then
16                         candidatePaths.append((path, missedScore, selectedCount))
17
18                   end if
19
20             end if
21
22       end foreach
23
24      if candidatePaths is not empty then
25             // Exploitation: Pick highest missed path
26             highestPath = max(candidatePaths, key=missedScore)
27             pathHistory [method][highestPath] += 1
28             selectedPaths[method].add(highestPath)
29             // Exploration: Pick least selected path
30             leastPath = min(candidatePaths, key=selectedCount)
31             if leastPath $\neq$ highestPath then
32                   pathHistory [method][leastPath] +=1
33                   selectedPaths[method].add(leastPath)
34
35             end if
36
37       end if
38
39 end foreach
return selectedPaths
Algorithm 2 Path Selection Strategy
In addition to the global data variables, the input to Algorithm 2 consists of the source file, the current test file, and the extracted paths for all methods under test (MUTs) in the source file, denoted as methodDict, obtained from Algorithm 1. The output consists of the selected paths for each MUT, which will be incorporated into the prompt for test generation.
[…]
For complex methods with high cyclomatic complexity, the large number of candidate execution paths poses a challenge, making it difficult for LLMs to process efficiently and potentially exceed the available context window. For instance, an untested method with a cyclomatic complexity of 20 could have 20 candidate paths, each representing a distinct execution flow.
To mitigate this, instead of providing all candidate paths in the prompt, we rank them and select two paths per method, based on exploitation and exploration (Lines 22). For exploitation, we choose the path with the highest missedScore to maximize code coverage improvement. For exploration, we select the path that has been visited the least in previous iterations, ensuring diversity in test generation. If the two selected paths are the same, we include it only once and update pathHistory accordingly. This approach ensures that LLMs have a clearer target and can focus more effectively on generating relevant test cases, without exceeding their context window.
## Source File
Here is the source file that you will be writing tests against.
{source_file_numbered}
## Test File
Here is the file that contains the existing tests.
{test_file}
## Third-party dependencies for test generation
Please use the following dependencies to generation tests
{test_dependencies}
{%- if failed_tests_section%}
{{ failed_tests_section }}
{endif %}
## Methods Under Test
Please generate test for {method_name_1} to cover the path {selected_path_for_method_1}
Please generate test for {method_name_N} to cover the path {selected_path_for_method_N}
Figure 4. Code Structure Aware Prompt Template
3.3. Program Analysis Guided Prompt Template
After selecting the paths, we construct a test generation prompt by inserting the selected paths into a structured prompt template. Our program analysis-guided prompt template, shown in Figure 4, consists of five sections: source file, test file, test dependencies, failed tests, and methods under test.
The source file section includes the complete source code with additional line numbers to facilitate mapping the selected path lines back to their corresponding locations in the code. The test file section contains the currently existing test cases. The test dependencies section specifies the required packages and their versions in the test scope, such as junit:4.13.2, to prevent LLMs from generating tests using incorrect package versions. The failed tests section includes test cases from the last iteration that failed, providing feedback to guide LLMs in refining test generation (subsection 3.4). Finally, the methods under test section includes the selected paths for each MUT, obtained using Algorithm 2. The paths follow the format illustrated in Figure 1, where each statement is mapped to its corresponding line number in the original source file, and conditional statements are annotated with their branch evaluation (true or false).
3.4. Iterative Test Generation Framework
Input: srcFileDir, testFileDir
1 testDeps = extractDependenciesForTestScope()
2 srcFile = readFile(srcFileDir)
3 if testFileDir does not exist then
4       testFile = initialTestClassSkeleton(testFileDir, testDeps)
5
6else
7       testFile = readFile(testFileDir)
8
9 end if
10 iter, iterNoIncrease = 0, 0
11 pathHistory = dict(), failedTestFeedback = []
12 methodDict = Algorithm 1 (srcFile)
13 maxCYC = getMaxComplexity(methodDict)
14 curCoverage = runTestsWithCoverage(testFile)
15
16while iter ¡ maxCYC and iterNoIncrease ¡ maxNoIncreaseLimit and curCoverage ¡ 100%  do
17       // build prompt for test generation
18       if curCoverage = 0% then
19             prompt = buildPrompt(srcFile, testFile, testDeps)
20
21      else
22             selectedPaths = Algorithm 2 (srcFile, testFile, methodDict, pathHistory)
23             prompt = buildPrompt(srcFile, testFile, testDeps, selectedPaths, failedTestFeedback)
24
25       end if
26
27      generatedTests = generateTestByPromptLLM(prompt)
28
29      // Validate generated tests
30       foreach test in generatedTests do
31             stdout = addNewTestAndRun(test, testFile)
32             if test is failed then
33                   errorMsg = parseErrorMessage(stdout)
34                   removeAddedTest(test, testFile)
35                   testsToRepair.add((test, errorMsg))
36
37             end if
38
39       end foreach
40       failedTestFeedback = repairAndValidate (testsToRepair, numOfAttempts)
41       newCoverage = runTestsWithCoverage(testFile)
42       if newCoverage ¿ curCoverage then
43             iterNoIncrease = 0
44
45      else
46             iterNoIncrease += 1
47
48       end if
49      curCoverage = newCoverage, iter += 1
50
51 end while
Algorithm 3 Iterative Test Generation Framework
In the previous subsections, we introduced the algorithms and components used to construct the hybrid program analysis-guided prompt for test generation, focusing on a single iteration. Now, we elaborate on the design of our iterative test generation workflow, which is presented in Algorithm 3 and explain how the various components integrate into the overall Panta framework.
Test Generation. Algorithm 3 begins by preparing the source file, target test file, and test dependencies (Lines 33). For example, if the project uses Maven, we can extract test dependencies using the command mvn dependency:list. We then read the source file as context from the directory $srcFileDir$. Our approach supports two scenarios: (1) augmenting existing tests, and (2) generating a new test file from scratch. We first check whether a test file exists at $testFileDir$. If a test file is available, we analyze and augment the existing test case; otherwise, we create a new test file based on the extracted test dependencies, initializing it with the test class declaration and necessary import statements.
We initialize workflow tracking variables (Lines 33) to monitor the stopping conditions. Specifically, we use iter to count the total number of iterations and iterNoIncrease to track the number of consecutive iterations without code coverage improvement. We hypothesize that if several consecutive iterations fail to enhance code coverage, it likely indicates convergence to a fixed value, making further execution unnecessary. Path selection data is stored in pathHistory, which is referenced by Algorithm 2 for selecting paths. Additionally, we maintain failedTestFeedback to store failed tests from the previous iteration, providing feedback to refine the current test generation process. Next, we perform program analysis to extract paths for MUTs using Algorithm 1 and obtain the initial code coverage information (Lines 33). We calculate the maximum cyclomatic complexity among all MUTs in the source file, denoted as maxCYC (Line 3). Since the number of paths for any MUT is at most maxCYC, we use this value to approximate the complexity of the source file and set it as the maximum number of iterations for test generation. This ensures that the iterative workflow runs for at most maxCYC iterations. Along with this iteration limit, the process terminates under three conditions (Line 3): (1) if the number of iterations reaches the limit maxCYC, (2) if the code coverage reaches 100%, or (3) if there are maxNoIncreaseLimit consecutive iterations without coverage improvement. These conditions help optimize resource usage and prevent unnecessary test generation attempts.
For each iteration, we determine the test generation strategy based on the current code coverage. If the coverage is 0%, indicating that no existing tests are available, we prompt LLMs to generate tests using only the source file, test file, and test dependencies, without providing the selected paths. This approach allows greater flexibility for LLMs to generate tests covering different parts of the source file (Lines 33). If the coverage is greater than 0%, we guide LLMs using our program analysis-guided prompt, which includes the selected paths to focus test generation on uncovered parts of the code (Lines 33).
Test Validation. After the LLM completes test generation, the generated tests are directed to the validation step (Lines 33). The tests are executed one by one. Passed tests are added to the test file, while failed tests, along with their error messages, are collected for the test repair process (Line 3). We prompt the LLM with the failed tests and their corresponding error messages, allowing it to attempt repairs for up to numOfAttempt times per failed test. We observed that if a test initially fails due to a compilation error, fixing the compilation error may reveal additional compilation errors or runtime errors. Therefore, a single repair attempt is often insufficient. To address this, we set the default value of numOfAttempt to 3. The repair process stops when either all failed tests are successfully fixed or the maximum number of repair attempts is reached. After the repair attempts, if any tests still fail, we store them along with their error messages in failedTestFeedback, which serves as feedback for the next round of test generation. Finally, we execute the current test file to obtain the updated coverage information and update the iter and iterNoIncrease variables (Lines 33). If none of the stop conditions have been met, the process continues with the next iteration.
4. Evaluation
We assess Panta to answer the following three research questions:
  • RQ1: How does the quality of tests generated by Panta compare to state-of-the-art test generation techniques?
  • RQ2: What is the individual contribution of each component of Panta to the overall quality of the generated tests?
  • RQ3: How do different LLMs affect the performance of Panta in test generation?
Table 1. Included subjects from Defects4J
Identifier Project #Class #MUTs $\textbf{CYC}_{max}$ $\textbf{CYC}_{med}$
Cli Cli-40f 2 31 11 1
Codec Codec-18f 7 78 31 1
Collections Collections-28f 5 218 32 2
Compress Compress-47f 9 65 22 3
Csv Csv-16f 3 72 25 1
Gson Gson-16f 4 75 37 2
JCore JacksonCore-26f 9 161 30 4
JDatabind JacksonDatabind-112f 9 371 22 2
JXml JacksonXml-5f 4 118 29 2
Jsoup Jsoup-93f 8 171 22 1
JxPath JxPath-22f 12 115 32 4
Lang Lang-4f 17 586 30 2
Math Math-2f 30 452 31 2
Time Time-13f 11 458 30 2
Total/Ave 130 2971 27 2
4.1. Experimental Setup
Subject Programs. We evaluate Panta using real-world subjects from Defects4J (Just et al., 2014). Specifically, we selected the latest fixed version of each subject program from Defects4J v2.0.1. To maintain consistency, we exclude three outdated projects—Chart, Mockito, and Closure—as they do not use Maven and rely on deprecated dependencies. The remaining 14 subjects are compatible with Java 8, Maven (v3.6.3), and JUnit (v4.13.2), and they integrate with JaCoCo (v0.8.11) (geeksforgeeks, 2025a) for code coverage measurement and Pitest (v1.17.0) (Coles et al., 2016) for mutation score evaluation.
To evaluate coverage for complex code, rather than generating tests for all classes, we selected non-abstract, public classes that contain at least one method with a cyclomatic complexity greater than 10 (McCabe, 1976; Sanusi et al., 2020). Table 1 presents the selected subjects. The column Project denotes the class name along with its version number from Defects4J. The column #Class represents the number of classes for which we generate tests, while #MUTs counts the total number of methods under test within the selected classes, for each project. In total, we evaluate 130 classes containing 2,971 methods under test (MUTs). Additionally, we report the maximum and median cyclomatic complexity values of the methods under test in the columns $\text{CYC}_{max}$, $\text{CYC}_{med}$, respectively.
Table 2. Comparing Panta with SymPrompt
Identifier #Class line coverage (%) branch coverage (%) pass rate (%) mutation score (%) HCC(#)
Panta SymPrompt Panta SymPrompt Panta SymPrompt Panta SymPrompt Panta SymPrompt
Cli 2 83.56 42.93 70.02 34.02 85.21 29.31 39.31 33.53 1 0
Codec 7 81.83 47.67 77.05 41.77 72.82 29.04 26.10 19.14 5 1
Collections 5 72.66 61.77 68.53 64.48 40.01 49.83 63.43 55.56 2 3
Compress 9 56.95 21.90 46.19 17.31 43.40 20.29 25.16 9.09 1 0
Csv 3 76.99 46.47 56.74 35.48 69.88 34.45 69.45 48.45 0 0
Gson 4 77.07 40.81 60.95 32.57 72.96 10.44 46.17 23.74 1 0
JCore 9 68.43 36.21 60.88 33.58 61.74 36.32 29.79 12.13 3 1
JDatabind 9 56.77 35.64 47.10 29.93 28.99 23.11 35.34 39.85 1 0
JXml 4 67.81 26.32 62.99 22.26 64.18 26.31 41.10 7.82 1 1
Jsoup 8 82.99 36.99 61.98 25.62 68.93 22.92 56.69 32.89 2 0
JxPath 12 56.80 24.40 49.05 22.48 50.97 14.08 45.97 20.73 4 1
Lang 17 71.18 57.71 64.13 52.18 72.27 51.39 39.01 53.31 5 2
Math 30 70.94 50.45 62.81 44.75 41.38 35.57 43.05 36.68 13 6
Time 11 80.57 58.66 70.34 47.41 71.86 45.05 52.45 45.33 6 1
Total/Ave 130 70.18 43.91 60.83 38.17 55.92 32.83 43.79 31.30 45 16
Evaluation Metrics. We adopt commonly used quality metrics in test generation literature (Wang et al., 2024a; Deljouyi et al., 2024; Ryan et al., 2024), including line coverage, branch coverage, and pass rate for RQ1 and RQ2. The pass rate is defined as the percentage of generated tests that successfully pass when added to the test suite. Since we evaluate at the class level rather than the method level, we measure line and branch coverage per class and report the average coverage across all targeted classes for each project. For RQ1 and RQ3, in addition to these standard metrics, we use two additional metrics: High Coverage Count (HCC) and mutation score. Based on industry standards (geeksforgeeks, 2025b; atlassian, 2025), typical code coverage targets range from 70% to 80%, e.g., Salesforce sets a 75% threshold for line coverage (salesforce, 2025). However, since line coverage is generally easier to achieve than branch coverage, we focus primarily on branch coverage. Therefore, we report the number of classes that achieve at least 75% branch coverage as a high coverage count, denoted as HCC. Additionally, we evaluate mutation score to assess test effectiveness, as coverage alone does not ensure fault detection if test assertions are non-existent or weak (Zhang and Mesbah, 2015).
LLM Selection. We select the latest open-source model, Meta Llama 3.3 70B (Llama, 2025), as our primary model for evaluation. For RQ3, to assess the impact of different models, we evaluate Panta on three additional models: one open-source model, Mistral Large 2 (Mistral, 2025) from Mistral AI, and two proprietary cost-efficient models—GPT-4o Mini from OpenAI (OpenAI, 2025) and Claude 3.5 Haiku from Anthropic (Anthropic, 2025).
Panta supports both augmenting existing test classes and generating new ones. To ensure a fair comparison, we evaluate Panta in the scenario where tests are generated for a new class without any existing test cases across all research questions.
Table 3. Ablation Study
Identifier #Class line coverage (%) branch coverage (%) pass rate (%)
baseline Panta_basic Panta_cov Panta baseline Panta_basic Panta_cov Panta baseline Panta_basic Panta_cov Panta
Cli 2 24.29 78.72 85.12 83.56 13.12 63.99 71.93 70.02 28.57 77.64 66.25 85.21
Codec 7 39.06 77.91 65.55 81.83 32.24 69.46 57.86 77.05 41.51 62.67 59.85 72.82
Collections 5 30.30 66.85 64.97 72.66 29.90 63.74 66.32 68.53 60.80 76.38 75.04 40.01
Compress 9 17.76 47.40 47.59 56.95 14.02 39.06 39.86 46.19 19.88 49.90 40.54 43.40
Csv 3 12.11 66.18 59.34 76.99 6.00 51.25 44.49 56.74 28.70 61.11 61.02 69.88
Gson 4 28.00 62.00 75.46 77.07 17.00 52.30 55.85 60.95 52.92 64.47 64.46 72.96
JCore 9 25.47 40.70 42.37 68.43 15.89 33.17 35.02 60.88 42.60 42.65 45.32 61.74
JDatabind 9 17.90 40.49 38.43 56.77 12.17 30.95 29.74 47.10 37.10 47.37 43.23 28.99
JXml 4 12.26 53.70 38.08 67.81 4.04 40.89 30.37 62.99 25.22 60.71 37.50 64.18
Jsoup 8 42.35 65.80 77.62 82.99 27.99 49.85 61.02 61.98 55.90 52.41 65.19 68.93
JxPath 12 30.77 53.28 52.35 56.80 23.89 45.41 45.42 49.05 42.08 59.60 47.51 50.97
Lang 17 34.47 60.90 58.72 71.18 29.85 53.74 51.34 64.13 65.31 67.29 64.92 72.27
Math 30 37.47 66.73 62.35 70.94 29.39 56.14 55.58 62.81 61.84 55.69 54.96 41.38
Time 11 55.47 71.42 68.84 80.57 40.96 59.40 57.84 70.34 86.71 65.77 66.53 71.86
$\text{CYC}{max}\leq 20$ 104 34.01 61.56 60.07 70.37 26.29 51.80 51.86 61.09 51.42 58.21 55.60 56.10
$\text{CYC}_{max}>20$ 26 27.38 55.49 52.68 69.42 19.17 46.00 42.93 59.81 56.55 59.15 56.66 55.16
Total/Ave 130 32.69 60.34 58.59 70.18 24.87 50.64 50.08 60.83 52.45 58.40 55.81 55.92
4.2. RQ1: Generated Test Quality Comparison
The latest test generation techniques relevant to Panta include HITS (Wang et al., 2024a) and SymPrompt (Ryan et al., 2024). HITS decomposes focal methods into slices and generates tests slice by slice, while SymPrompt extracts path constraints as a list of branch conditions for each method under test and generates tests path by path. Neither approach leverages dynamic code coverage data and both operate statically without guidance on where additional tests are needed. Consequently, HITS and SymPrompt are limited to generating new test classes rather than augmenting existing ones. To differentiate Panta with them, Panta adopts an iterative workflow, guiding the LLM to generate tests for selected paths based on both dynamic and static program analysis information. Our approach works for both scenarios to augment existing tests and add new test classes.
However, neither HITS nor SymPrompt has publicly available implementations. HITS relies on an LLM to generate slices instead of traditional program analysis, and the paper (Wang et al., 2024a) provides limited implementation details, making replication challenging. In contrast, SymPrompt provides sufficient details to replicate its approach. Therefore, to compare it with Panta, we implemented SymPrompt’s technique following the information that is available in their paper (Ryan et al., 2024).
We present the results in Table 2. We highlight the best values in bold to enhance visualization in the result tables. In terms of line and branch coverage, Panta consistently outperforms SymPrompt across all projects.
On average, Panta achieves 26.3% higher line coverage and 22.7% higher branch coverage than SymPrompt. Regarding high coverage count (HCC), 45 out of 130 classes achieve branch coverage above 75%, compared to only 16 for SymPrompt. The pass rate of SymPrompt is, on average, 23.1% lower than Panta. Beyond the absence of a test repair phase, this can be attributed to its limited prompt context. Specifically, SymPrompt provides only branch conditions rather than the full execution path, making it more challenging for the LLM to infer correct test inputs. In terms of mutation score, Panta achieves a higher average score (43.8%) than SymPrompt (31.3%), aligning with its superior code coverage results.
There are a few scenarios in which SymPrompt performs better. In the Collections project, SymPrompt achieves a higher pass rate than Panta, while in JDatabind and Lang, its mutation score exceeds that of Panta. Upon analysis, these scenarios occur in large utility classes containing large number of methods. For example, MapUtils in Collections has 71 methods. Since Panta generates tests at the class level, it selects paths for each MUT that have not been fully covered in each iteration. For large classes, this results in long prompts, increasing the difficulty for the LLM to generate complete tests. In contrast, SymPrompt operates at the method level, allowing it to focus on smaller contexts, making it more effective in these specific cases.
Overall, across all evaluation metrics—line coverage, branch coverage, pass rate, HCC, and mutation score—Panta consistently outperforms SymPrompt.
4.3. RQ2: Ablation study
To assess the contributions of each component of Panta, we design different variants that provide the LLM with varying levels of information for test generation. We evaluate Panta in a scenario where tests are generated for a new class without any existing test cases. As shown in Algorithm 3, the essential inputs include the source file (srcFile), test dependencies (testDeps), and a new test file (testFile), initialized with the test class declaration and necessary import statements.
Table 4. Comparison among different models
Identifier line coverage (%) branch coverage (%) pass rate (%) mutation score (%) HCC(#)
Llama Mistral GPT Claude Llama Mistral GPT Claude Llama Mistral GPT Claude Llama Mistral GPT Claude Llama Mistral GPT Claude
Cli 83.56 58.04 75.79 97.61 70.02 39.38 54.41 82.33 85.21 33.52 66.73 86.30 39.31 33.53 47.98 74.57 1 0 0 2
Codec 81.83 73.20 55.24 86.20 77.05 68.00 49.45 80.82 72.82 41.40 43.72 69.48 26.10 36.37 25.35 48.14 5 4 1 5
Collections 72.66 78.64 60.73 61.10 68.53 77.61 58.13 62.95 40.01 51.59 53.64 82.73 63.43 70.89 62.87 62.87 2 4 2 3
Compress 56.95 31.14 37.52 74.83 46.19 27.44 29.13 70.08 43.40 16.99 23.67 61.61 25.16 11.82 13.98 30.05 1 0 0 4
Csv 76.99 69.17 46.39 62.08 56.74 58.61 28.77 46.89 69.88 33.16 24.23 55.13 69.45 67.78 52.98 68.02 0 0 0 0
Gson 77.07 48.32 73.32 79.33 60.95 37.67 56.48 61.24 72.96 20.02 34.58 62.03 46.17 25.29 47.99 53.05 1 1 0 1
JCore 68.43 45.73 44.75 65.90 60.88 40.25 39.27 61.04 61.74 31.65 35.72 69.16 29.79 13.32 13.61 25.48 3 2 1 4
JDatabind 56.77 45.96 52.50 52.45 47.10 42.99 46.70 49.70 28.99 23.22 49.83 62.81 35.34 37.14 35.83 38.12 1 2 1 2
JXml 67.81 41.88 34.59 34.02 62.99 38.66 32.41 38.17 64.18 31.23 31.88 25.11 41.10 17.64 13.04 15.95 1 1 1 1
Jsoup 82.99 60.00 68.25 80.22 61.98 45.75 52.87 65.62 68.93 38.82 48.68 75.29 56.69 42.70 45.45 59.65 2 1 1 3
JxPath 56.80 50.85 32.66 65.85 49.05 42.17 29.20 57.78 50.97 29.07 29.04 52.02 45.97 43.38 29.90 50.48 4 1 1 4
Lang 71.18 65.30 65.00 75.61 64.13 58.85 58.64 71.57 72.27 39.63 56.24 82.44 39.01 32.43 33.05 39.42 5 5 4 9
Math 70.94 56.04 55.02 81.27 62.81 49.62 49.65 72.87 41.38 40.77 51.01 73.03 43.05 28.30 30.56 44.86 13 8 8 16
Time 80.57 70.99 64.26 79.81 70.34 63.61 54.25 74.15 71.86 45.37 47.05 64.81 52.45 51.85 35.46 60.24 6 5 2 8
Total/Ave 70.18 56.60 54.05 73.20 60.83 49.85 46.77 66.52 55.92 35.48 44.39 68.16 43.79 36.60 34.86 47.92 45 34 22 62
To establish a baseline, we prompt the LLM with only this basic information for a single test generation pass, without leveraging multiple iterations, test repair, or feedback, which are key features of our iterative framework. To assess the impact of our iterative framework, we introduce Panta_basic, a variant that generates prompts consistently across all iterations using only the basic information, without incorporating any program analysis data. Additionally, we evaluate Panta_cov to determine whether code coverage information alone can enhance test coverage. The Panta_cov variant extends the basic information by incorporating code coverage data, providing the LLM with the locations of uncovered lines and branches in the current class as line numbers mapped to the original source. However, it does not include path selection details on how to cover these uncovered elements.
The results comparing the baseline, Panta_basic, Panta_cov, and Panta are presented in Table 3. In addition to reporting the average line/branch coverage and pass rate for each project, we classify the 130 evaluated classes based on their maximum cyclomatic complexity ($\text{CYC}_{max}$), using two groups: $\text{CYC}{max}\leq 20$ and $\text{CYC}_{max}>20$. A cyclomatic complexity greater than 20 is considered an indicator of highly complex programs (Sanusi et al., 2020).
For line and branch coverage, we observe a substantial improvement when comparing Panta_basic to the baseline, with line coverage increasing by 27.65% and branch coverage by 25.77%. This demonstrates that our iterative framework significantly enhances overall code coverage. While Panta_cov achieves similar coverage, its average coverage is slightly lower than that of Panta_basic. A Paired T-test (Marascuilo and Serlin, 1988) confirms that the difference between them is not statistically significant across all subjects. Notably, for the Cli project, which has lower complexity ($\text{CYC}_{max}=11$), Panta_cov achieves the highest line and branch coverage. This suggests that for less complex subjects, code coverage information alone can be beneficial, but for more complex projects, it is insufficient to drive meaningful improvements in coverage.
By integrating code coverage-driven path selection (Algorithm 2), Panta further improves line and branch coverage by approximately 10% compared to the other two Panta variants. When analyzing coverage across complexity categories, we observe a notable drop of 6% to 9% for highly complex subjects ($\text{CYC}_{max}>20$) in the baseline, Panta_basic, and Panta_cov variants. In contrast, Panta maintains nearly consistent coverage across both complexity levels, with only a 1 percentage point difference. These findings demonstrate that our iterative framework, combined with hybrid program analysis leveraging both static and dynamic information, effectively enhances code coverage, particularly for complex subjects.
The pass rate is primarily determined by the LLM’s ability to generate correct code. Since all variants use the same LLM (Llama 3.3 70B), the pass rate remains relatively stable. Panta_basic achieves a 6% higher pass rate than the baseline, attributed to the test repair phase incorporated in our framework. However, after integrating program analysis information into the prompt, the overall pass rate is slightly lower than Panta_basic, with only a 3.5% improvement over the baseline. This trend holds for both Panta_cov and Panta. The reason is that prompts in Panta_cov and Panta guide the LLM to target more complex and hard-to-cover execution paths, making it more challenging to generate correct test cases. In contrast, Panta_basic offers greater flexibility, allowing the LLM to decide where to add tests. As a result, the LLM tends to avoid complex paths, leading to a higher pass rate by focusing on easier-to-cover cases.
4.4. RQ3: Impact of different models
In addition to Llama 3.3 70B, we evaluate Panta on three other models, as introduced in Section 4.1. Table 4 presents the results across four models: Llama 3.3 70B (Llama), Mistral Large 2 (Mistral), GPT-4o Mini (GPT), and Claude 3.5 Haiku (Claude) for each evaluation metric.
For code coverage metrics, Claude performs the best, achieving 73.2% line coverage, 66.5% branch coverage, and 62 out of 130 classes surpassing 75% branch coverage. Llama follows, with 70.2% line coverage, 60.8% branch coverage, and 45 classes exceeding the 75% threshold. GPT performs the worst in achieving high coverage. Thus, the ranking for code coverage is Claude ¿ Llama ¿ Mistral ¿ GPT. The same ranking applies to mutation score, as Claude also achieves the highest score, followed by Llama, Mistral, and GPT. For pass rate, which reflects the model’s ability to generate correct code, Claude again performs the best (68.2%), with Llama as the second highest (55.9%). However, unlike the coverage metrics, GPT achieves a higher pass rate than Mistral, resulting in a ranking of Claude ¿ Llama ¿ GPT ¿ Mistral for pass rate.
Overall, Claude is the best-performing model, but as a proprietary model, it has the highest API pricing among the four models. For a more cost-effective alternative, Llama achieves results comparable to Claude while being open-source, making it a more budget-friendly option.
5. Discussion
Limitations. Since our approach focuses on analyzing one class at a time, its effectiveness diminishes for classes that rely on inheritance. These classes often override parent class methods, resulting in the absence of explicit method implementations, which reduces the effectiveness of control flow analysis. In the future, we plan to extend our work to incorporate cross-class context to address this limitation. Regarding test repair, a common limitation across different LLMs is their inability to effectively fix runtime errors, such as AssertionError caused by failed test assertions. In contrast, LLMs perform better in resolving compilation errors. This indicates that LLMs lack the capability to fully understand project-specific execution logic, as they struggle to infer runtime behaviors and expected outcomes. Regarding test effectiveness, we observe that the average mutation score across different LLMs consistently falls below 50%, indicating that less than half of the introduced mutants are detected. This suggests that LLM-generated tests still have significant room for improvement in enhancing fault detection capability.
Application. The tests generated by Panta operate under the assumption that they are created within a regression testing setting, where the correctness of the underlying focal method implementation is presumed. As Panta supports both augmenting existing test classes and generating new ones, it can be integrated with any test generation tool to further improve code coverage in their generated test suites.
6. Threats To Validity
Internal Validity. The effectiveness of Panta depends on the capabilities of the LLM used, and results may vary if more advanced models are introduced in the future. To mitigate this, we evaluate Panta using four recently released LLMs, including two open-source and two proprietary models. Another potential issue is the relationship between code coverage and test effectiveness, as high coverage does not necessarily imply meaningful assertions. To address this, we include mutation score in our evaluation to assess fault-detection capability alongside traditional coverage metrics. Additionally, Panta integrates Comex (Das et al., 2023) to generate control flow graphs for path extraction. Any bugs or limitations in Comex could potentially influence our results. To mitigate this risk, we verify the extracted paths during the implementation phase to ensure their correctness before they are used for test generation.
External Validity. Our current implementation and evaluation focus on Java, which may impact the generalizability of our findings. However, both LLM-based test generation and program analysis techniques used in Panta are language-agnostic. The integration of dynamic code coverage analysis and static control flow analysis can be extended to support multiple programming languages, provided that equivalent tooling is available.
Reproducibility. As this work is still under review, we will make Panta’s implementation and dataset publicly available soon.
7. Related Work
Automated unit test generation has been a longstanding research topic, with approaches historically falling into two main categories: search-based and learning-based. Popular search-based approaches, such as EvoSuite (Fraser and Arcuri, 2011, 2013) for Java and Pynguin (Lukasczyk and Fraser, 2022; Lukasczyk et al., 2020) for Python, utilize evolutionary algorithms to generate high-coverage unit tests. However, these approaches often suffer from generating test cases that are difficult for developers to understand and may lack meaningful assertions (Rao et al., 2023; Deljouyi et al., 2024). On the other hand, learning-based approaches leverage deep learning techniques to improve test case quality. For instance, ATLAS (Watson et al., 2020) employs a neural machine translation model to automatically generate meaningful assertion statements for test methods, while AthenaTest (Tufano et al., 2020) uses a transformer-based model trained on developer-written test cases. These approaches produce more readable and accurate test cases, addressing key limitations of search-based methods.
With the rapid evolution of Large Language Models (LLMs), their ability to generate code has significantly improved, leading to their widespread adoption for automated test generation related tasks. Meta’s TestGen-LLM (Alshahwan et al., 2024) employs LLMs to automatically enhance existing human-written tests. Libro (Kang et al., 2023) generates test cases directly from bug reports. Cedar (Nashid et al., 2023) introduces an automated retrieval-based demonstration selection method to improve test assertion generation. CodaMosa(Lemieux et al., 2023) enhances the search-based test generation tool Pynguin(Lukasczyk and Fraser, 2022; Lukasczyk et al., 2020) by incorporating LLMs to overcome coverage plateaus. UTGen (Deljouyi et al., 2024) integrates search-based software testing with LLMs to improve the understandability of unit tests generated by EvoSuite. ChatUniTest (Chen et al., 2024) follows a generation-validation-repair mechanism to refine LLM-generated tests. ASTER (Pan et al., 2025) describes a generic pipeline applicable to multiple programming languages and complex software systems that require environment mocking. Empirical studies have also been conducted to evaluate the test generation capabilities of different LLMs across various programming languages, including JavaScript (Schäfer et al., 2023), Java (Yang et al., 2024; Siddiq et al., 2024), and Python (Wang et al., 2024b).
However, the empirical studies (Siddiq et al., 2024; Schäfer et al., 2023; El Haji, 2023; Wang et al., 2024b) highlight a persistent challenge: LLMs struggle to generate unit tests for complex methods and often fail to achieve sufficient branch coverage, particularly in real-world projects. To address this limitation, recent research has explored strategies to provide LLMs with additional program-related information beyond just the focal class or focal method, enhancing their ability to generate tests for complex code.
HITS (Wang et al., 2024a) prompts LLMs to decompose focal methods into slices and generate tests for each slice. However, these ”slices” are not derived from traditional program slicing techniques (Weiser, 1984) but instead represent arbitrary method fragments generated by LLMs.
SymPrompt (Ryan et al., 2024) employs symbolic execution to extract path constraints from each method under test. It then reduces the extensive set of extracted paths to a manageable subset of constraints and generates tests incrementally, path by path. In contrast, Panta leverages a path approximation algorithm that directly computes a linearly independent set of execution paths from the control flow graph. Moreover, while SymPrompt provides only branch conditions, Panta delivers full execution paths, enabling the LLM to more accurately infer the correct test inputs.
Additionally, both SymPrompt and HITS provide static prompts containing method fragments (slices or paths) but lack guidance on where additional tests are needed, restricting them to generating new test classes rather than augmenting existing ones. CoverUp (Pizzorno and Berger, 2024), on the other hand, provides code coverage information to indicate which parts of the code remain uncovered but does not guide the LLM on how to cover those uncovered segments effectively.
Unlike existing techniques, Panta adopts an iterative workflow that leverages hybrid program analysis to guide the LLM. It integrates dynamic code coverage analysis and static control flow analysis to systematically extract and rank uncovered paths. These paths then help the LLM in reasoning about where and how to generate tests. Also, unlike previous methods, our approach effectively supports both augmenting existing tests and generating new test classes when no tests exist for a class/method.
8. Conclusion
This paper introduced Panta, a novel technique for automated unit test generation that seeks to emulate how human developers create tests. It addresses the challenge of achieving high branch coverage in complex code, where large language models often fall short. Panta employs an iterative approach, combining static control flow analysis to identify execution paths with dynamic code coverage analysis to pinpoint under-tested areas. This hybrid program analysis guides an LLM to generate targeted test cases, which are then validated and repaired, using feedback to refine subsequent test generation efforts. Our evaluations on complex classes from the Defects4J (Just et al., 2014) benchmark show Panta significantly outperforms existing methods in line and branch coverage, achieving 26% higher line coverage and 23% higher branch coverage compared to the state-of-the-art, demonstrating the effectiveness of its iterative, program analysis-guided strategy for improving test thoroughness.
Looking ahead, we plan to extend Panta to support multiple programming languages, broadening its applicability across diverse software projects. Additionally, we aim to incorporate cross-class context to further reduce hallucination and improve the accuracy of generated test cases.